OpenStudy (sparklestaraa):

A wooded area in a state park has a mixture of different types of trees. There are 800 pine trees and 50 oak trees. The number of pine trees is decreasing at a rate of 5% per year. The number of oak trees is increasing at a rate of 15% per year. If these trends continue:

OpenStudy (sparklestaraa):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

write what you have so far

OpenStudy (sparklestaraa):

f(x)=800-5?

OpenStudy (sparklestaraa):

*50

OpenStudy (sparklestaraa):

thats all i have :( because I'm always confused with functions they don't make much sense to me

jimthompson5910 (jim_thompson5910):

how many pine trees do we start with?

OpenStudy (sparklestaraa):

800

jimthompson5910 (jim_thompson5910):

we have the number of pine trees decreasing by 5% each year so we subtract 5% of 800 from 800 800 - (5% of 800) = 800 - 0.05*800 = 800*(1-0.05) = 800*0.95 basically we take 95% of 800 to get the next value

OpenStudy (sparklestaraa):

i have these functions right y = 800(.95)^x y = 50(1.15)^x

jimthompson5910 (jim_thompson5910):

good, you nailed both

OpenStudy (sparklestaraa):

now for B?

OpenStudy (sparklestaraa):

During what year in the future will the park have approximately the same number of pine and oak trees?

OpenStudy (sparklestaraa):

The answer is 16 and 14.5 years right? but idk how to get to that answer

jimthompson5910 (jim_thompson5910):

one sec

jimthompson5910 (jim_thompson5910):

probably the best way is to graph both y = 800(.95)^x and y = 50(1.15)^x and look for when they cross

jimthompson5910 (jim_thompson5910):

do you have a graphing calculator?

OpenStudy (sparklestaraa):

no i never had one :(

OpenStudy (sparklestaraa):

this is b right?

jimthompson5910 (jim_thompson5910):

ok you can use something like desmos https://www.desmos.com/calculator

jimthompson5910 (jim_thompson5910):

it's a free online graphing calculator

OpenStudy (sparklestaraa):

omg xD i have tried that countless times and it always shows that little red line in some random place

OpenStudy (sparklestaraa):

but let me try

jimthompson5910 (jim_thompson5910):

what did you type in?

OpenStudy (sparklestaraa):

y=800(.95)^x

jimthompson5910 (jim_thompson5910):

zoom out until you see a red line

jimthompson5910 (jim_thompson5910):

it should be a curve actually

OpenStudy (sparklestaraa):

k i see it now so the second one

OpenStudy (sparklestaraa):

do i put it a different color and how do i do that

OpenStudy (sparklestaraa):

i got that too sorry asking needless questions ok :D I'm done with the graph

jimthompson5910 (jim_thompson5910):

it should automatically put it in a different color for you

jimthompson5910 (jim_thompson5910):

nope not needless, they're good questions

jimthompson5910 (jim_thompson5910):

you can move the graph around by clicking on it

jimthompson5910 (jim_thompson5910):

try to see where they cross

jimthompson5910 (jim_thompson5910):

you can click the point of intersection and desmos will display (or try to) the point of intersection

OpenStudy (sparklestaraa):

(14.51,380.03)

jimthompson5910 (jim_thompson5910):

I'm getting the same

jimthompson5910 (jim_thompson5910):

the x represents how much time has gone by the y represents the amount of trees

jimthompson5910 (jim_thompson5910):

both values are approximate

OpenStudy (sparklestaraa):

so this is B?

jimthompson5910 (jim_thompson5910):

it's used to answer B and C

OpenStudy (sparklestaraa):

so b is not 16 and 14.5

jimthompson5910 (jim_thompson5910):

how are you getting "16 and 14.5"

OpenStudy (sparklestaraa):

Im not i just figured i did both Decrease and Increase formulas but I'm wrong

OpenStudy (sparklestaraa):

so heres all of what i wrote :D 4. A. y = 800(.95)^x y = 50(1.15)^x B. 14.5 years C. 380 tress approximately in 14.5 years

OpenStudy (sparklestaraa):

oh i have to graph those 2 functions

jimthompson5910 (jim_thompson5910):

for C, they just want the number of trees

jimthompson5910 (jim_thompson5910):

but yeah you have the right idea

OpenStudy (sparklestaraa):

:D ok so don't write 14.5

jimthompson5910 (jim_thompson5910):

yeah I think 380 works fine

OpenStudy (sparklestaraa):

ok well Honestly thank you so so so so much :) if i have anymore problems i'll send you a message

jimthompson5910 (jim_thompson5910):

you're welcome